The lenghts of the sides of a triangle are 25cm,60cm,and 65 cm.If the longest side of a similar trinagle measures 39cm,what is the lenght of the shortest side of this triangle
PLEASE HELP!!!!!!!!!!!!!!!!!!

Answer by stanbon(48535)

The lenghts of the sides of a triangle are 25cm,60cm,and 65 cm.If the longest side of a similar trinagle measures 39cm,what is the lenght of the shortest side of this triangle
-----------------------
shortest / shortest = longest / longest
-----
x/25 = 65/39
x = 25(65/39)
x = 41 2/3 cm
================
Cheers,
Stan H.